# Extragalactic Stellar Tidal Streams: Observations meet Simulation

### v1

@@ Abstract
Context. According to the well established hierarchical framework for galaxy evolution, galaxies grow through mergers with other galaxies and the $\Lambda$CDM cosmological model predicts that the stellar halos of galaxies are rich in remnants from minor mergers. The Stellar Streams Legacy Survey has provided a first release of a catalogue with a statistically significant sample of stellar streams in the Local Universe that can be used to study minor mergers and test the cosmological models.
Aims. The main objective is to compare the results of the observations of stellar tidal streams with the predictions of state-of-the-art cosmological simulations regarding the formation of stellar streams up to a redshift z ¡ 0.02, according to the $\Lambda$CDM model.
Methods. We use the predictions of the cosmological simulations Copernicus Complexio, TNG50 of the IllustrisTNG project and Auriga to generate 225 mock-images of nearby halos at a distance of 70 Mpc, and search for stellar streams. We compare the obtained stream frequency and characteristics with those obtained from the Stellar Streams Legacy Survey.
Results. We find good agreement between the results of analysing real images from the Dark Energy Survey and mock-images from cosmological simulations. We obtained predictions for the detection rate of stellar streams to a surface brightness limit of 35 mag arcsec-2
Conclusions. The cosmological simulations predict that for a surface brightness limit of 32 mag arcsec-2 a frequency of almost 70% in the detection of streams around galaxies can be achieved.
@@ 1 Introduction
According to the well established hierarchical framework for galaxy evolution, galaxies grow through mergers with other galaxies. These mergers can be major mergers, when the merging galaxies are of similar stellar mass; a mass ratio ¿ $1/3$ is a generally accepted threshold, see e.g. Newberg & Carlin (2016), and minor mergers, when the host galaxy accretes a dwarf galaxy in its halo. The Lambda Cold Dark Matter ($\Lambda$CDM) cosmological model predicts that the stellar halos of galaxies are rich in remnants from minor mergers that, in the Local Universe, on the basis of observations and simulations, are expected to be more frequent than major mergers (Guo & White, 2008; Jackson et al., 2022).
@@
Figure 1: Examples of stellar stream images from the Stellar Streams Legacy Survey (Martínez-Delgado et al., 2023; Miro-Carretero et al., 2024). The cutout angular sizes vary between 3x3 and 10x10 arcmin. Credits: G. Donatiello/D.Martínez-Delgado.
@@
The detection and study of extragalactic stellar tidal streams contributes to augment the stream census, mostly built so far from streams detected in the Milky Way and Local Volume (Belokurov et al., 2006; Martínez-Delgado et al., 2010; Hood et al., 2018; Shipp et al., 2018; Ferguson et al., 2022; Li et al., 2022), and helps us to contrast their frequency and characteristics with the predictions of the $\Lambda$CDM model on a statistically sound basis (the Local Volume is considered here to be a spherical region with a radius of 11 Mpc around the Milky Way or up to a radial velocity of redshift of z ¡ 0.002). This motivates the search for streams beyond the Local Volume, and up to a distance for which surveys are available with the required depth (for the most part streams are expected to be fainter than $\sim$ 30 mag arcsec-2, e.g. Johnston et al., 2008). Due to their low surface brightness much less data can be gathered for each individual stream in a survey of distant hosts, in comparison to the streams in our galaxy or in the Local Volume. In particular, star-by-star photometric and kinematic measurements are not possible for streams at these distances.
@@
A number of relevant surveys of tidal features beyond the Local Volume have been reported in the last decade: Duc et al. (2015); Morales et al. (2018); Bílek et al. (2020); Sola et al. (2022); Martínez-Delgado et al. (2023); Miró-Carretero et al. (2023); Giri et al. (2023); Rutherford et al. (2024); Skryabina et al. (2024); Miro-Carretero et al. (2024). Skryabina et al. (2024) present the results of visual inspection of a sample of 838 edge-on galaxies using images from three surveys: SDSS Strip-82, Subaru HSC and DESI (DECals, MzLS, BASS). This study, like our present work, was motivated to construct a deep photometric sample and obtain better statistics of tidal structures in the Local Universe in order to compare with cosmological simulations. The definition of tidal features used in that study includes also disc deformations and tidal tails, typical of major mergers. Their results will be discussed further in Section 6.
@@
The use of cosmological simulations to study tidal features has also increased significantly in the recent times (Martínez-Delgado, 2019; Mancillas et al., 2019; Martin et al., 2022; Vera-Casanova et al., 2022; Valenzuela & Remus, 2024; Khalid et al., 2024) Relevant work on stream detectability using mock-images from cosmological simulations is reported in Vera-Casanova et al. (2022). The authors have inspected surface brightness maps generated from 30 Auriga project simulations (Grand et al., 2017) of Milky Way-like galaxies looking for the brightest streams. They report that no streams have been detected in images with a surface brightness limit brighter than 25 mag arcsec-2. Their stream detection frequency increases significantly between 28 and 29 mag arcsec-2. They find a correlation between infall time and infall mass of the stream progenitors, such that more massive progenitors tend to be accreted at later times.
@@
In this work we use the term stellar tidal streams to refer to the remnants of minor mergers, in line with the nomenclature used in Martínez-Delgado et al. (2023)111great circles, hereinafter referred to as circles, are streams that result from satellites along mildly eccentric orbits, with an arc-like shape, sometimes featuring complete loops around the host, but in most cases (in our sample) seen as covering only a small part of a loop; umbrellas, structures often appearing on both sides of the host galaxy, displaying an elongated shaft ending in the form of a shell (sometimes only the shells are visible) resulting from satellites that were on more eccentric, radial orbits; giant plumes, hereinafter referred to as plumes, structures appearing to shoot out of the host, generally for quite a long distance. Our work focuses on stellar tidal streams in the Local Universe up to a distance of 100 Mpc (redshift $z<0.02$). We consider as stellar tidal streams only those low surface brightness (LSB) features that are of an accreted origin, whatever their apparent morphology (shells, circles, plumes etc.); as we will discuss later in the paper, the apparent morphology is strongly dependent on the line of sight of the observation. We can broadly characterise stellar tidal streams as LSB structures in the halo of galaxies, at distances between $\sim$ 20 and 120 kpc from the host centre and with surface brightness fainter than $\sim$ 25 mag arcsec-2. Stellar tidal streams are thus a particular case of LSB structures and are significantly (several mag arcsec-2) fainter than tidal tails, another type of LSB feature resulting primarily from major mergers (Toomre & Toomre, 1972).
The results from observations of the Dark energy Survey (DES) presented in Miro-Carretero et al. (2024) allow for a direct, quantitative comparison of the abundance and characteristics of stellar tidal streams in the Local Universe with the predictions from state-of-the-art cosmological simulations based on the $\Lambda$CDM paradigm. In particular, we can compare statistics derived from the observed stream population (for example, the number of stream detections at at given surface brightness limit, or the distribution of photometric observables for detected streams) with those predicted by cosmological simulations. To do this, we obtain predictions of stream formation from three cosmological simulations: Copernicus Complexio (COCO, Hellwing et al., 2016), Illustris TNG50 (Pillepich et al., 2019; Nelson et al., 2019, 2019a) and Auriga (Grand et al., 2017).
@@
In this paper, we compare the stellar streams in a sample of galaxies observed by the DES survey with streams in mock images derived from the simulations listed above, for a matched sample of hosts. We compare the detection frequency and photometric characteristics measured in both samples and discuss the results. In Section 2 we introduce the main characteristics of the cosmological simulations. The selection of the halos from the simulations to be analysed is presented in Section 3. Section 4 is devoted to the process of generating mock images. In Section 5 we present the predicted detectability of streams at different surface brightness limits. The results of the comparison are discussed in Section 6 and the summary, conclusions and outlook are given in Section 7.
@@ 2.1 Copernicus Complexio
The Copernicus Complexio (COCO. Hellwing et al., 2016) is a $\Lambda$CDM cosmological $N$-body simulation post-processed with a semi-analytic galaxy formation model and the ‘Stellar Tags in N-Body Galaxy Simulations’ (STINGS) particle tagging technique (Cooper et al., 2010, 2017). COCO provides both high mass resolution and an approximate analogue of the Local Volume (a high-resolution spherical region of radius $\sim$25 Mpc, with density slightly lower than the cosmic mean, embedded in a lower-resolution box of 100 Mpc/side). The Galform semi-analytic model of Lacey et al. (2016) was used to predict the evolution of the baryonic component in each dark matter halo. This model is calibrated to a range of low and high-redshift observables, including optical and near-IR luminosity functions, the HI mass function, and the relationship between the masses of bulges mass and central supermassive black holes. The 6-dimensional phase space of each single-age stellar population formed in the model is mapped to an individual subset of dark matter particles using the STINGS technique. These models will be presented in detail in a future publication (Cooper et al. in prep.). The specific characteristics of the COCO simulations are listed in 1.
2.2 Illustris TNG50
IllustrisTNG is a suite of large volume, cosmological, magnetohydrodynamical simulations run with the moving-mesh code AREPO (Springel 2010). The Illustris TNG50 (Pillepich et al., 2019) simulations characteristics are listed in Table 1:
The IllustrisTNG50 simulation includes a comprehensive model for galaxy formation physics, which is able to realistically follow the formation and evolution of galaxies across cosmic time (Weinberger et al., 2017; Pillepich et al., 2018). Each IllustrisTNG50 simulation self-consistently solves the coupled evolution of dark matter, cosmic gas, luminous stars, and supermassive blackholes from a starting redshift of $z=127$ to the present day, $z=0$. We select the TNG50-1 simulation run and the snapshots at $z<0.02$.
@@ 2.3 Auriga
The Auriga simulations (Grand et al., 2017, 2024) are a set of cosmological zoom-in magnetohydrodynamical simulations also carried out with the AREPO code. Auriga re-simulates at higher resolution a sample of halos selected by the EAGLE project (Schaye et al., 2015). From the available simulations in the Auriga portal [^2] https://wwwmpa.mpa-garching.mpg.de/auriga/data.html we use all 30 halos in the Original/4 series.
Figure 2: Stellar mass versus B-band absolute magnitude from the Spitzer S4G catalogue. The dashed line represents the empirical correlation in equation 1.
@@ 3 Halo Selection
The histogram in Figure 3 shows the stellar mass distribution of the DES sample and of the selected halos in the COCO, IllustrisTNG50 and Auriga simulations. The average stellar mass of the DES sample is $\log_{10}M_{\star}/\mathrm{M_{\odot}}$ = 10.00 and the standard deviation $\sigma=0.38$. While the average of the distributions is not the same, there is sufficient overlap between the DES sample stellar mass range and the stellar mass range of the simulated halos selected for the comparison. Figure 4 shows the halo mass distribution of the selected halos in the COCO, IllustrisTNG50 and Auriga simulations. Here all the simulations overlap and the average values are close to one another. In this stellar mass range, COCO galaxies occupy a broader range of halo masses than IllustrisTNG50 galaxies.
@@ 4 Mock Images
To compare the predictions of cosmological simulations with the observations from the DES sample, we have generated mock images from snapshots of the simulations described in Section 2 in the redshift range $0.020<z<$ 0.02. The continuous stellar mass density field in all these simulations is represented by discrete tracers, called star particles [^4] In the COCO simulation, stellar mass is associated with a subset of the collisionless particles in post-processing, rather than an independent particle species in the original simulation, but the principle is the same from the point of view of our analysis; more details are given in Cooper et al. (2017).. The stellar mass associated with each star particle corresponds to a stellar population with a single age and metallicity. As in any $N$-body realization of a density field, each particle notionally corresponds to an irregular volume of phase space centered on the location of the particle. We have applied the following transformations to the properties of the star particles from the simulation snapshots in order to recover the observables that can be identified in real images:
@@
For the expansion (smoothing) step, we distribute the luminosity (flux) of each particle over the image pixels by convolution with a Gaussian kernel. The scale of this kernel, $h$, which we refer to as the smoothing scale, is set to a parameterized multiple to the root-mean-square average distance to the 16th-nearest star particle neighbour, $h_{16}$. It therefore adapts to the local star particle density. The logic for setting $h$ is broadly similar to that used to determine the kernel scale in a smoothed particle hydrodynamics calculation. However, the smoothing in our case only serves to interpolate between the original particles and does not have any physical significance. It is therefore somewhat arbitrary; our method reflects a balance between smoothing sufficiently to reduce the visual impression of a discrete particle distribution, while preserving the small-scale features.
@@
Figure 7: Comparison of smoothing methods. Images generated from the surface brightness maps, by transforming the surface brightness to counts, and adding background noise to emulate a surface brightness limit of 29 mag arcsec-2. Left and centre images correspond to a smoothing length equal to the distances to the 16th nearest neighbour multiplied by a factor of $0.3$ (left) and a factor of $0.6$ (center). The right panel shows the image produced with a different smoothing length, considering the distance to the 5th neighbour.
We use the mock images to assess the detectability of streams over a range of surface brightness limits. To follow the same process of stellar stream detection by visual inspection as with real images, we have transformed the surface brightness maps into counts images. Then we have added background noise to the images, which depends on the type of analysis to be carried out.
@@
Figure 8 shows an example of a COCO galaxy to which we have added the sky background of a real DES image with a surface brightness limit of $28.65\,\mathrm{mag\,arcsec^{-2}}$. We compare this to the same mock image combined with a simple Gaussian noise background equivalent to the same surface brightness limit. A stream with a umbrella /shell morphology can be clearly appreciated in the image with the real DES background and its appearance is consistent with that in the image with an artificial background. Adding a synthetic background to the mock image does not seem to impact significantly the detection of streams by means of visual inspection with respect to the real background in this surface brightness limit regime. This method is very efficient to emulate different levels of surface brightness limit and we adopt it in this work to assess the detectability to those levels currently not achievable in available surveys. However we acknowledge that for much fainter surface brightness limits, this method provides only an approximation as the confusion of sources and possibly cirri will become much more significant and the difference between the synthetic background and the real background will increase.
Figure 8: Example of a COCO simulation halo stacked on a real DES sky background (right), compared with the mock images with artificial background corresponding to the same surface brightness limit of 28.65 [mag $arcsec^{-}2$] (left).

### v2

@@ Abstract
Context. According to the well established hierarchical framework for galaxy evolution, galaxies grow through mergers with other galaxies and the $\Lambda$CDM cosmological model predicts that the stellar halos of massive galaxies are rich in remnants from minor mergers. The Stellar Streams Legacy Survey (SSLS) has provided a first release of a catalogue with a statistically significant sample of stellar streams in the local Universe, detected in deep images from DESI Legacy Surveys and the Dark Energy Survey (DES).
Aims. The main objective is to compare observations of stellar tidal streams from the SSLS catalogue with predictions from state-of-the-art cosmological simulations regarding their abundance, up to a redshift z ¡ 0.02, according to the $\Lambda$CDM model.
Methods. In particular, we use the outcome of the cosmological simulations Copernicus Complexio, TNG50 of the IllustrisTNG project, and Auriga to generate mock images of nearby halos and search for stellar streams. We compare the stream frequency and characteristics found in these images, as well as the results of a photometric analysis of the simulations data, with DES observations.
Results. We find generally good agreement between the real images and the simulated ones regarding frequency and photometry of streams, while the stream morphology is somewhat different between observations and simulations, and between simulations themselves. By varying the sky background of the synthetic images to emulate different surface brightness limit levels, we also obtain predictions for the detection rate of stellar tidal streams up to a surface brightness limit of 35 mag arcsec-2.
Conclusions. The cosmological simulations predict that with an instrument such as the one used in the DES, it would be necessary to reach a surface brightness limit of 32 mag arcsec-2 in the $r$-band to achieve a frequency of up to $\sim$ 70% in the detection of stellar tidal streams around galaxies in the redshift range considered here.
@@ 1 Introduction
According to the well established hierarchical framework for galaxy evolution, galaxies grow through mergers with other galaxies. These mergers can be major mergers, when the merging galaxies are of similar stellar mass; a mass ratio ¿ $1/3$ is a generally accepted threshold, see e.g. Newberg & Carlin (2016), and minor mergers, when the host galaxy accretes a dwarf galaxy in its halo. The Lambda Cold Dark Matter ($\Lambda$CDM) cosmological model predicts that the stellar halos of massive galaxies ($\log_{10}M_{\star}/M_{\odot}\gtrsim 9$) are rich in remnants from minor mergers that, in the local Universe, on the basis of observations and simulations, are expected to be more frequent than major mergers (Guo & White, 2008; Jackson et al., 2022).
@@
Figure 1: Examples of stellar stream images from the Stellar Streams Legacy Survey (Martínez-Delgado, 2024). The cutout angular sizes vary between 3x3 and 10x10 arcmin.
@@
The detection and study of extragalactic stellar tidal streams contributes to augment the stream census, mostly built so far from streams detected in the Milky Way (MW) and Local Volume (Belokurov et al., 2006; Martínez-Delgado et al., 2010; Hood et al., 2018; Shipp et al., 2018; Ferguson et al., 2022; Li et al., 2022), and helps us to contrast their frequency and characteristics with the predictions of the $\Lambda$CDM model on a statistically sound basis (the Local Volume is considered here to be a spherical region with a radius of 11 Mpc around the MW or up to a radial velocity of redshift of z ¡ 0.002). This motivates the search for streams beyond the Local Volume, and up to a distance for which surveys are available with the required depth (for the most part streams are expected to be fainter than $\sim$ 30 mag arcsec-2, e.g. Johnston et al., 2008). Due to their low surface brightness much less data can be gathered for each individual stream in a survey of distant hosts, in comparison to the streams in our galaxy or in the Local Volume. In particular, star-by-star photometric and kinematic measurements are usually not possible for streams at these distances.
@@
A number of relevant surveys of tidal features beyond the Local Volume have been reported in the last decade: Duc et al. (2015); Morales et al. (2018); Bílek et al. (2020); Sola et al. (2022); Martínez-Delgado et al. (2023); Miró-Carretero et al. (2023); Giri et al. (2023); Rutherford et al. (2024); Skryabina et al. (2024); Miro-Carretero et al. (2024). Skryabina et al. (2024) present the results of visual inspection of a sample of 838 edge-on galaxies using images from three surveys: SDSS Strip-82, Subaru HSC and DESI (DECals, MzLS, BASS). This study, as our present work, was motivated to construct a deep photometric sample and obtain better statistics of tidal structures in the local Universe in order to compare with cosmological simulations. The definition of tidal features used in that study includes also disc deformations and tidal tails, typical of major mergers. Their results will be discussed further in Section 7.
@@
The use of cosmological simulations to study tidal features has also increased significantly in the recent times (Martínez-Delgado, 2019; Mancillas et al., 2019; Martin et al., 2022; Vera-Casanova et al., 2022; Valenzuela & Remus, 2024; Khalid et al., 2024) Relevant work on stream detectability using mock-images from cosmological simulations is reported in Vera-Casanova et al. (2022). The authors have inspected surface brightness maps generated from 30 Auriga project simulations (Grand et al., 2017) of MW-like galaxies looking for the brightest streams. They report that no streams have been detected in images with a surface brightness limit brighter than 25 mag arcsec-2. Their stream detection frequency increases significantly between 28 and 29 mag arcsec-2. They find a correlation between infall time and infall mass of the stream progenitors, such that more massive progenitors tend to be accreted at later times.
@@
In this work we use the term stellar tidal streams to refer to the remnants of minor mergers, in line with the nomenclature used in Martínez-Delgado et al. (2023)111great circles, hereinafter referred to as circles, are streams that result from satellites along mildly eccentric orbits, with an arc-like shape, sometimes featuring complete loops around the host, but in most cases (in our sample) seen as covering only a small part of a loop; umbrellas, structures often appearing on both sides of the host galaxy, displaying an elongated shaft ending in the form of a shell (sometimes only the shells are visible) resulting from satellites that were on more eccentric, radial orbits; giant plumes, hereinafter referred to as plumes, structures appearing to shoot out of the host, generally for quite a long distance. Our work focuses on stellar tidal streams in the local Universe up to a distance of 100 Mpc (redshift $z<0.02$). We consider as stellar tidal streams only those low surface brightness (LSB) features that are of an accreted origin, whatever their apparent morphology (shells, circles, plumes etc.); as we will discuss later in the paper, the apparent morphology is strongly dependent on the line of sight of the observation. We can broadly characterise stellar tidal streams as LSB structures in the halo of galaxies, at distances between $\sim$ 20 and 120 kpc from the host centre and with surface brightness fainter than $\sim$ 25 mag arcsec-2 for the most part. Stellar tidal streams are thus a particular case of LSB structures and are significantly (several mag arcsec-2) fainter than tidal tails, another type of LSB feature resulting primarily from major mergers (Toomre & Toomre, 1972).
The results from observations of the Dark energy Survey (DES) presented in Miro-Carretero et al. (2024) allow for a direct, quantitative comparison of the abundance and characteristics of stellar tidal streams in the local Universe with the predictions from state-of-the-art cosmological simulations based on the $\Lambda$CDM paradigm. In particular, we can compare statistics derived from the observed stream population (for example, the number of stream detections at at given surface brightness limit, or the distribution of photometric observables for detected streams) with those predicted by cosmological simulations, as well as the comparison of the measured photometry. To do this, we obtain predictions of stream formation from three cosmological simulations: Copernicus Complexio (COCO, Hellwing et al., 2016), TNG50 (Pillepich et al., 2019; Nelson et al., 2019a, b) and Auriga (Grand et al., 2017).
@@
In this paper, we compare the stellar streams in a sample of galaxies observed by the DES survey with streams in mock images derived from the simulations listed above, for a matched sample of hosts. We compare the detection frequency and photometric characteristics measured in both samples and discuss the results. In Section 2 we introduce the main characteristics of the cosmological simulations. The selection of the halos from the simulations to be analysed is presented in Section 3. Section 4 is devoted to the process of generating mock images. In Section 5 we present the predicted detectability of streams at different surface brightness limits. In Section 6 we present the results of the photometry measurements in the mock images. The results of the comparison are discussed in Section 7 and the summary, conclusions and outlook are given in Section 8.
@@ 2.1 Copernicus Complexio
The Copernicus Complexio (COCO. Hellwing et al., 2016) is a $\Lambda$CDM cosmological $N$-body simulation post-processed with a semi-analytic galaxy formation model and the ‘Stellar Tags in N-Body Galaxy Simulations’ (STINGS) particle tagging technique (Cooper et al., 2010, 2017). COCO provides both high mass resolution and an approximate analogue of the Local Volume (a high-resolution spherical region of radius $\sim$25 Mpc, with density slightly lower than the cosmic mean, embedded in a lower-resolution box of 100 Mpc/side). The Galform semi-analytic model of Lacey et al. (2016) was used to predict the evolution of the baryonic component in each dark matter halo. This model is calibrated to a range of low and high-redshift observables, including optical and near-IR luminosity functions, the HI mass function, and the relationship between the masses of bulges mass and central supermassive black holes. The 6-dimensional phase space of each single-age stellar population formed in the model is mapped to an individual subset of dark matter particles using the STINGS technique. These models will be presented in detail in a future publication (Cooper et al. in prep.). The specific characteristics of the COCO simulations are listed in Table 1.
2.2 TNG50
IllustrisTNG is a suite of large volume, cosmological, magnetohydrodynamical simulations run with the moving-mesh code AREPO (Springel, 2010) and with publicly available data (Nelson et al., 2019a). One of such runs, the highest resolution, is known as TNG50-1, or simply TNG50 (Pillepich et al., 2019; Nelson et al., 2019b) and is used in this work. Its key characteristics are listed in Table 1.
The TNG50 simulation includes a comprehensive model for galaxy formation physics, which is able to realistically follow the formation and evolution of galaxies across cosmic time (Weinberger et al., 2017; Pillepich et al., 2018). TNG50 self-consistently solves the coupled evolution of dark matter, cosmic gas, luminous stars, and supermassive black-holes from a starting redshift of $z=127$ to the present day, $z=0$. In this work, we focus on the snapshots at $z<0.02$.
@@ 2.3 Auriga
The Auriga simulations (Grand et al., 2017, 2024) are a set of cosmological zoom-in magnetohydrodynamical simulations also carried out with the AREPO code. Auriga re-simulates at higher resolution a sample of halos selected by the EAGLE project (Schaye et al., 2015) and implements models for black hole (BH) accretion and feedback, stellar feedback, stellar evolution, chemical evolution, metallicity dependent cooling, star formation and magnetic fields, as reported in Grand et al. (2017). From the available simulations in the Auriga portal [^2] https://wwwmpa.mpa-garching.mpg.de/auriga/data.html we use all 30 halos in the Original/4 series.
Figure 2: Stellar mass versus B-band absolute magnitude from the Spitzer S4G catalogue. The red line represents the empirical correlation in equation 1.
@@
Figure 5: Stellar mass vs halo mass for the selected sample of galaxies in the COCO, TNG50 and Auriga simulations compared with the empirical correlation by Moster et al. (2018) $\pm$ $2\sigma$. The horizontal lines indicate the DES galaxy sample stellar mass distribution: mean value (dotted line); mean value $-1\sigma$ (lower dashed line) and mean value $+2\sigma$ (upper dashed line). The yellow star represents the MW.
@@ 3 Halo Selection
The histogram in Figure 3 shows the stellar mass distribution of the DES sample and of the selected halos in the COCO, TNG50 and Auriga simulations. The average stellar mass of the DES sample is $\log_{10}M_{\star}/\mathrm{M_{\odot}}$ = 10.00 and the standard deviation $\sigma=0.38$. While the average of the distributions is not the same, there is sufficient overlap between the DES sample stellar mass range and the stellar mass range of the simulated halos selected for the comparison. Figure 4 shows the halo mass distribution of the selected halos in the COCO, TNG50 and Auriga simulations. Here all the simulations overlap and the average values are close to one another. In this stellar mass range, COCO galaxies occupy a broader range of halo masses than TNG50 galaxies.
Figure 5 shows the stellar mass versus the halo mass for the selected sample of galaxies in the COCO, TNG50 and Auriga simulations compared with the empirical correlation by Moster et al. (2018) $\pm$ $2\sigma$. The horizontal lines indicate the mean value, mean value $-1\sigma$ and mean value $+2\sigma$ for the DES galaxy sample stellar mass distribution. While TNG50 galaxies are within Moster’s correlation $\pm$ $2\sigma$, some Auriga galaxies seem to be slightly above the $=2\sigma$ correlation range. COCO shows a larger scatter of stellar mass in the region above $\log_{10}M_{halo}/M_{\odot}$ and has a rather sparse sample of galaxies around the MW region. These COCO simulation characteristics are discussed in depth in Cooper et al. in prep.
@@ 4 Mock Images
To compare the predictions of cosmological simulations with the observations from the DES sample, we have generated mock images from snapshots of the simulations described in Section 2 in the redshift range $0.020<z<$ 0.02. The continuous stellar mass density field in all these simulations is represented by discrete tracers, called star particles [^4] In the COCO simulation, stellar mass is associated with a subset of the collisionless particles in post-processing, rather than an independent particle species in the original simulation, but the principle is the same from the point of view of our analysis; more details are given in Cooper et al. (2017).. The stellar mass associated with each star particle corresponds to a stellar population with a single age and metallicity. As in any $N$-body realization of a density field, each particle notionally corresponds to an irregular volume of phase space centred on the location of the particle. We have applied the following transformations to the properties of the star particles from the simulation snapshots in order to recover the observables that can be identified in real images:
@@
For the expansion (smoothing) step, we distribute the luminosity (flux) of each particle over the image pixels by convolution with a Gaussian kernel. The scale of this kernel, $h$, which we refer to as the smoothing scale, is set to a parametrized multiple to the root-mean-square average distance to the 16th-nearest star particle neighbour, $h_{16}$. It therefore adapts to the local star particle density. The logic for setting $h$ is broadly similar to that used to determine the kernel scale in a smoothed particle hydrodynamics calculation. However, the smoothing in our case only serves to interpolate between the original particles and does not have any physical significance. It is therefore somewhat arbitrary; our method reflects a balance between smoothing sufficiently to reduce the visual impression of a discrete particle distribution, while preserving the small-scale features.
@@
Figure 8: Comparison of smoothing methods. Images generated from the surface brightness maps, by transforming the surface brightness to counts, and adding background noise to emulate a surface brightness limit of 29 mag arcsec-2. Left and centre images correspond to a smoothing length equal to the distances to the 16th nearest neighbour multiplied by a factor of $0.3$ (left) and a factor of $0.6$ (centre). The right panel shows the image produced with a different smoothing length, considering the distance to the 5th neighbour.
We use the mock images i) to assess the detectability of streams over a range of surface brightness limits and ii) to measure the photometry of the streams that would be detectable at the surface brightness limit of the DES observations. To follow the same process of stellar stream detection by visual inspection as with real images, we have transformed the surface brightness maps into counts images. Then we have added background noise to the images, which depends on the type of analysis to be carried out.
@@
To measure photometric properties of the mock images and compare the results with those of the DES sample, we add a more realistic sky background to the mock images. We first extract the sky background from selected real DES sample images, removing the central galaxy and replacing it by real sky background from an area of the same image without significant point sources. The image from which we extracted this fiducial background image was selected according to the following criteria: i) the central galaxy should be edge-on, in order to minimise its area of influence on the image; ii) the image should have no stream detection, in order not to interfere with the synthetic halo image; and iii) the image should have an $r$-band surface brightness limit representative of the DES sample. The selected image had a surface brightness limit of 28.65 mag arcsec-2. The fiducial sky background extracted from this image, as above, was then superimposed onto the mock images generated directly from the surface brightness maps, that have only a central galaxy and possibly other accompanying galaxies but no sky background, creating an image with a synthetic central galaxy and a real DES background.
We use Gnuastro to replace the central galaxy and its surroundings with a section of the same image extracted from a region without bright sources. The size and orientation of the ellipsoidal mask is such that covers the area of influence of the central galaxy in its surroundings, i.e. it reaches up to the point where the surface brightness profile flattens.
Figure 9 shows examples of TNG50 and COCO mock images with streams around the central galaxies. Panels B show images to which we have added the sky background of a real DES image with a surface brightness limit of $28.65\,\mathrm{mag\,arcsec^{-2}}$. We compare these images to those in panels A, to which we have added an artificial Gaussian sky background equivalent to the same surface brightness limit. The streams, though faint, can be clearly appreciated in the images with real DES background and, although they are more obvious in the image with artificial background, the their appearance is consistent in both images. Consequently, adding a synthetic background to the mock images does not seem to impact significantly the detection of streams by means of visual inspection with respect to the real background in this surface brightness limit regime. This method is very efficient to emulate different levels of surface brightness limit and we adopt it in this work to assess the detectability of streams at those levels currently not achieved in available surveys. However we acknowledge that for much fainter surface brightness limits, this method provides only an approximation as the confusion of sources and possibly cirri will become much more significant and the difference between the synthetic background and the real background will increase.
Figure 9: Comparison of mock images with simulated and real DES sky background: examples of a TNG50 (left) and a COCO (right) simulated galaxies with streams stacked with a real DES sky background (panels B), and with artificial flat background (panels A) both having the same surface brightness limit of 28.65 [mag arcsec-2].